OpenStudy (anonymous):

a slide 5.7 m long makes an agle of 32 degrees with the ground. how hight is the slide above the ground?

OpenStudy (amistre64):

sin(32) = h/5.7 h = 5.7(sin(32))

OpenStudy (anonymous):

thanks!